# Quantum DeepONet: Neural operators accelerated by quantum computing

### v1

@@ 2.1.4 Tomography for extracting classical output
If the quantum circuit include quantum noise, error mitigation methods could be introduced during tomography based on the system’s properties. We apply the same method in Ref. [46], where only unary measurement outcomes are kept and all the other non-unary outcomes are discarded. This post-selection technique is a benefit of unary encoding. The effect of this method will be demonstrated in Section 4.3.2.
@@ 2.1.5 Summary and remarks
In conclusion, the structure of a complete quantum layer is shown in Fig. 1. The number of qubits needed is $1+\max(m,n)$ for $\mathbf{W}\in\mathbb{R}^{m\times n}$, in which the top $1$ qubit is an ancillary qubit for tomography purpose, while other qubits are used to store information and perform operations. Sequentially, we implement data loading, pyramidal circuit and tomography, and thus complete the matrix multiplication in quantum space.
@@
Quantum layers can accelerate the feedforward pass, achieving a complexity of $O(n/\delta^{2})$. Here, $\delta$ is the threshold for the tomography error. The complexities of other components of quantum layers are shown in Table 1.
@@ 2.2 Quantum orthogonal neural network
Figure 2: Architecture of quantum DeepONet. DeepONet consists of two subnetworks: the branch net and the trunk net. In quantum DeepONet, we replace these with QOrthoNN, which is composed of several quantum layers arranged sequentially. The nonlinear operations are performed on classical computers.
@@ 2.3 Quantum DeepONet
A DeepONet includes a branch net and trunk net, each with an equivalent number of output neuron, denoted by $p$. The branch and trunk nets can adopt arbitrary architectures, like fully connected neural network (FNN), convolutional neural network (CNN), recurrent neural network (RNN), and residual neural network (ResNet). A diagrammatic representation of DeepONet is illustrated in the center of Fig. 2. The branch network receives the input function evaluated at a discrete set of points $\{z_{1},z_{2},\ldots,z_{q}\}$, represented by $[v(z_{1}),v(z_{2}),\ldots,v(z_{q})]$. The trunk net is fed with the location $\xi$ at which the output function is evaluated, which can include both time and space coordinates. The outputs of the branch and trunk networks are denoted by $[b_{1}(v),b_{2}(v),\ldots,b_{p}(v)]$ and $[t_{1}(\xi),t_{2}(\xi),\ldots,t_{p}(\xi)]$. Thus, the final output of DeepONet is the sum of the dot product of the branch and trunk network outputs and a bias $b_{0}\in\mathbb{R}$, y expressed as
@@ 3 Ideal quantum simulation results
To demonstrate the efficacy of our method, we first use QOrthoNN to approximate certain functions (Section 3.1). Subsequently, we move to the application of quantum DeepONet on learning ODE and PDE problems, including the antiderivative operator (Section 3.2), advection equation (Section 3.3), and Burgers’ equation (Section 3.4). Finally, we test QPI-DeepONet using antiderivative operator and Poisson’s equation (Section 3.5).
We implement the classical training by using the library DeepXDE [8]. After classical training on OrthoNN, we extract the weights and biases and construct a quantum version incorporating quantum layers, applying Qiskit [54] for quantum simulation. It is important to note that, in this section, we adopt an idealized scenario during quantum simulation. This approach exclude any quantum and statistical noise, aiming to assess the theoretical accuracy and performance of the quantum model. The hyperparameters of neural networks and $L^{2}$ relative error of different examples are summarized in Table 3. The code of all examples are published in GitHub (https://github.com/lu-group/quantum-deeponet).
Table 3: The default parameters and test $L^{2}$ relative error for different examples of quantum DeepONet. For quantum DeepONet, the first number in the “Depth” column is the depth of branch net, and the second number is the depth of trunk net. The same for the “Activation” column. The “Error” column presents results for both classical training and ideal simulation, as they are identical for all examples.
@@ 3.1 Function approximation
and approximate it using OrthoNN. We choose 80 points for training and 100 points for testing, where $x$ is uniformly sampled in $[-1,1]$. Specially, for this example, we use tanh activation function to circumvent the “dying ReLU” problem [55], which is particularly relevant here given the small width of the network.
For this function 1, we can achieve a small $L^{2}$ relative error of $0.149\%$ for testing set after training classically. Following classical training, we construct QOrthoNN using the pyramidal quantum circuit, based on the classically training parameter. The ideal quantum simulation yields an error identical to classical training: $0.149\%$. Essentially, OrthoNN and QOrthoNN are the same neural network, differing only in their prediction methods—one is executed on a classical computer, while the other is run on a quantum simulator. The results for true function, OrthoNN and QOrthoNN are plotted in Fig. 3A, where the three lines align closely with each other.
Figure 3: Quantum simulation result of function predictions. The black, red and blue lines represent the reference solution, classical prediction of OrthoNN, and ideal quantum simulation result of QOrthoNN, respectively. (A) Results for $f(x)=1/(1+25x^{2})$. (B) Results for $f(x)=\sum_{k=1}^{4}\sin(kx)$.
@@ 3.2 Antiderivative operator
Next we examine quantum DeepONet by starting with an antiderivate operator:
@@
where $k_{l}(x_{i},x_{j})=\exp(-\frac{d(x_{i},x_{j})^{2}}{2l^{2}})$ denotes the radial basis function (RBF) kernel. In this context, $d(\cdot,\cdot)$ is the Euclidean distance between two points and $l$ represents the length scale of the kernel, which modulates the smoothness of the generated function. Specifically, an increase of the value of $l$ leads to a smoother generated function. Therefore, we can adjust $l$ depending on our desired level of function’s complexity.
In this example, we explore two scenarios with different length scales: $l=1.0$ and $l=0.5$, corresponding to different size of the network during training. We achieved small errors of $0.49\%$ and $0.84\%$, respectively in theses two scenarios for both OrthoNN and ideal simulation of QOrthoNN (see Table 3).
@@ 3.3 Advection Equation
$\frac{\partial u}{\partial t}+\frac{\partial u}{\partial x}=0,\quad x\in[0,1],~{}t\in[0,1],$
@@
$k(x_{i},x_{j})=\exp{\Big{(}-\frac{2\sin^{2}{(\pi d(x_{i},x_{j})/p)}}{l^{2}}\Big{)}}.$
Here, $p$ is the periodicity of the kernel and is set to $1$. We choose $l=1.5$ and derive the ground truth using the analytical solution $u(x,t)=u_{0}(x-t)$. For branch inputs $u_{0}(x)$, 20 sensors are uniformly placed (see one example in Fig. 4A left). Regarding the trunk inputs, we employ a grid of $50\times 50$ points, covering the range of $x$ and $t$. We implement the ResNet [56] architecture in both branch and trunk nets, which has a formulation of $\mathbf{x}^{\prime}=\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})+\mathbf{x}$ for each layer. This approach effectively mitigate the issue of gradient vanishing during training. The final test error of classical prediction reaches $2.25\%$. Ideal simulation of quantum DeepONet yeilds the same error: $2.25\%$. Fig. 4A provides an example of illustrating the ground truth, predictions of quantum DeepONet and the absolute error between them.
@@ 3.4 Burgers’ Equation
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},\quad x\in[0,1],~{}t\in[0,1],$
with initial condition $u_{0}(x)$ and periodic boundary condition, where $\nu=0.05$ is the viscosity. We aim to learn the mapping from $u_{0}(x)$ to the solution $u(x,t)$. Additionally, recognizing the periodic nature of the output function $u(x,t)$, instead of directly input $\xi$ for trunk net, we expand it to $[\xi,\cos(2\pi\xi),\sin(2\pi\xi),\cos(4\pi\xi),\sin(4\pi\xi)]$. Other neural network settings are the same as Section 3.3, except for the depth and width. The classical test relative error is $1.38\%$, and the ideal quantum simulation yields the same result: $1.38\%$. An example of the ground truth and prediction of quantum DeepONet is shown in Fig. 4B.
@@ 3.5 Quantum physics-informed DeepONet
In this section, we further show that our quantum DeepONet can also be trained using physics-informed loss term (Eq. (3)), without labeled data. We choose antiderivative euqation in Eq. (4) for comparison with data driven case in Section 3.2. Additionally, 1D Poisson’s equation
@@
with zero Dirichlet boundary condition is also considered for demonstration. The branch inputs in both cases are the $v(x)$ in equations, which is generated by GRF with RBF kernel. To facilitate training and keep PDE residual within reasonable range, in Poisson’s equation, we multiply generated GRF with a factor of 10 and take the enlarged function as input sample. The boundary condition is hard constrained using corresponding neural network architecture as mentioned in Section 2.5. Zero coordinate shift algorithm [57] is utilized to reduce GPU memory consumption and training time. During the training, the number of input samples is 10000 with batch size = 2000. Adam optimizing is used with $2\times 10^{5}$ iteration. The PDE residual is evaluated at 100 uniformly distributed points in $[0,1]$. During the training, we employed PCA with original dimension 100. The training results are shown in Table. 4. Ideal quantum simulation results also agree well with classical training results for all of these examples, shown “$L^{2}$ relative error” column.
We also conducted experiments without any dimension reduction techniques. For antiderivative with initial condition $l=1$, using branch and trunk net with depth of 3 and width of 10, the test $L^{2}$ relative error is $4.05\%$. For comparison, using the same hyperparameters with PCA, which projects original 100 dimensions down to 10, resulted in an error of $0.76\%$. We believe this difference is due to the critical dependency of QPI-DeepONet on the sampling of input sensors. PCA enables us to incoporate more information within limited input dimension. The limitations of current quantum devices compel us to use narrower neural networks, leading to sparse sampling of the branch input. As derivatives are taken with respect to the inputs, QPI-DeepONet is more sensitive to the input data.
@@ 4 Effects of noise
Quantum noise is a major obstacle for the practicality of a quantum algorithm in the noisy intermediate-scale quantum (NISQ) era. It emerges from various sources, including the imperfect implementation of quantum operators, undesired environmental or qubit interactions, and erroneous state preparation or measurement. During the execution of a quantum circuit, the accumulated errors produced by the noise can destroy any information we intend to obtain. Meanwhile, even in a fault-tolerant scenario, the inaccuracy resulting from the finite number of measurements still affects the error level and complexity of the neural network, making it unavoidable in the discussion of the feasibility of our work on near-term quantum computers. Thus, in Sections 4.1 and 4.2, we first provide a theoretical analysis of the effects of finite-sampling noise on the single RBS gate and tomography outputs and a well-known noise channel, depolarizing noise, respectively. Then, we demonstrate our noisy simulation results of quantum DeepONet under both types of noise, as well as a more comprehensive noise model emulating a real IBM quantum computer in Section 4.3.
@@ 4.1 Finite-sampling noise in tomography
In the tomography step (Section. 2.1.4), the probabilities ${\rm Pr}[0,\mathbf{e}_{j}]$ and ${\rm Pr}[1,\mathbf{e}_{j}]$, for $j\in\{1,...,r\}$, are estimated from the frequencies of measurement outcomes, where $n$ is the input vector dimension, $m$ is the output vector dimension, and $r=\max(m,n)$. This introduces an additional error on the estimation of output vector $\mathbf{y}$, caused by the finite number of measurements (shots). Let $q^{(0,j)}$ be the probability of measuring |$\ket{0,\mathbf{e}_{j}}$ in tomography layer, and $\hat{q}^{(0,j)}$ the estimated value from $N_{shot}$ shots. To estimate the size of the finite-sampling error, we first calculate the standard deviation of $\hat{q}^{(0,j)}$.
@@
and $S^{(0,j)}=\sum_{k=1}^{N_{shot}}Z^{(0,j)}_{k}$ is a Binomial random variable. Thus, we have
@@
Since the variance of $S^{(0,j)}$ is $\mathrm{Var}\left[S^{(0,j)}\right]=N_{shot}q^{(0,j)}\left(1-q^{(0,j)}\right)$, we have
$\mathrm{Var}\left[\hat{q}^{(0,j)}\right]=\mathrm{Var}\left[\frac{S^{(0,j)}}{N_{shot}}\right]=\frac{q^{(0,j)}\left(1-q^{(0,j)}\right)}{N_{shot}}.$ (5)
Similarly, define $\hat{q}^{(1,j)}$ as the probability of measuring |$\ket{1,\mathbf{e}_{j}}$ in tomography layer and let $\hat{q}^{(1,j)}$ be its estimate from $N_{shot}$ shots. We can obtain the variance of $\hat{q}^{(1,j)}$ similar to Eq. (5). It follows that
$\mathrm{Cov}\left[\hat{q}^{(0,j)},\hat{q}^{(1,j)}\right]$ $=\frac{1}{N^{2}_{shot}}\sum_{k=1}^{N_{shot}}\sum_{l=1}^{N_{shot}}\mathrm{E}\left[Z^{(0,j)}_{k}Z^{(1,j)}_{l}\right]-\mathrm{E}\left[\frac{S^{(0,j)}}{N_{shot}}\right]\mathrm{E}\left[\frac{S^{(1,j)}}{N_{shot}}\right]$
$=\frac{1}{N^{2}_{shot}}N^{2}_{shot}q^{(0,j)}q^{(1,j)}-q^{(0,j)}q^{(1,j)}$
$=0.$
With Eq. (1), the standard deviation of the estimated $y_{j}$ is
$\mathrm{Std}\left[y_{i}\right]$ $=\mathrm{Std}\left[\sqrt{r}({\rm Pr}[0,\mathbf{e}_{i}]-{\rm Pr}[1,\mathbf{e}_{j}])\right]$
$=\sqrt{r}\sqrt{\mathrm{Var}\left[\hat{q}^{(0,j)}\right]+\mathrm{Var}\left[\hat{q}^{(1,j)}\right]-2\mathrm{Cov}\left[\hat{q}^{(0,j)},\hat{q}^{(1,j)}\right]}$
$=\frac{\sqrt{r}}{\sqrt{N_{shot}}}\sqrt{q^{(0,j)}\left(1-q^{(0,j)}\right)+q^{(1,j)}\left(1-q^{(1,j)}\right)}$
$\propto\frac{\sqrt{r}}{\sqrt{N_{shot}}},$ (6)
where $q^{(0,j)},q^{(1,j)}\in[0,1]$. In conclusion, the finite-sampling error on the estimation of output vector $\mathbf{y}\in\mathbb{R}^{r}$ is proportional to $N_{shot}^{-0.5}$ when $r$ is relatively small compare to $N_{shot}$.

### v2

@@ 2.1.5 Error mitigation
If the quantum circuit includes quantum noise, error mitigation methods could be introduced during tomography based on the system’s properties. We apply the same method as in Ref. [46], where only unary measurement outcomes are kept and all the other non-unary outcomes are discarded. For example, if the measurement results in $0010$, it is a valid result to keep; if the measurement yields $0110$, then this result is ignored because it is not a unary string. This post-selection technique is a benefit of unary encoding. The improvement on the $L^{2}$ error from this method, along with the percentages of kept and discarded shots, will be demonstrated in Figs. 8B, C, and D in Section 4.3.2. In general, it provides a significant decrease in $L^{2}$ error across different levels of depolarizing error and finite-sampling error.
@@ 2.1.6 Summary and remarks
In conclusion, the structure of a complete quantum layer is shown in Fig. 1. The number of qubits needed is $1+\max(m,n)$ for $\mathbf{W}\in\mathbb{R}^{m\times n}$, in which the top $1$ qubit is an ancillary qubit for tomography purposes, while other qubits are used to store information and perform operations. Sequentially, we implement data loading, pyramidal circuit, and tomography, and thus complete the matrix multiplication in quantum space.
@@
Quantum layers can accelerate the feedforward pass, achieving a complexity of $O(n/\delta^{2})$. Here, $\delta$ is the threshold for the tomography error. This complexity is the result of the full $\mathcal{O}(n)$-depth circuit in Fig. 1 with $\mathcal{O}(1/\delta^{2})$ number of measurements in the data extraction. We will explicitly discuss the statistical error scaling in the data extraction in Section 4.1. The complexities of all the components of quantum layers are shown in Table 1. The depth of a single-layer circuit in Fig. 1 has at most $3n+\mathcal{O}(1)$ RBS gates. This is counted by first considering the bigger pyramidal group of RBS gates with depth at most $12n+$ 1 by combining the component “Load $\mathbf{x}$”, “Pyramid $\mathbf{W}$”, and “Load $1/\sqrt{r}$”. Then, the right-most “Load $1/\sqrt{r}$” has depth $n-1$ and there are some extra $\mathcal{O}(1)$ depths from the remaining gates related to the ancillary qubit. Providing the transpilation of the RBS gate in Appendix A, the circuit in Fig. 1 maintains an $\mathcal{O}(n)$ depth in terms of basis gates. Fig. 2 illustrates this linear relation with and without gate optimization from Qiskit.
@@
Figure 2: Circuit depth growth for a circuit of the single all-to-all quantum layer in Fig. 1. The basis gate set is $\{ECR,RZ,SX,X\}$. Two levels of gate optimization with Qiskit transpiler are considered: level 00 means no optimization and level $2$ includes qubit-layout optimization, inverse cancellation, 1-qubit gate optimization, and commutative cancellation.
@@ 2.2 Quantum orthogonal neural network
Figure 3: Architecture of quantum DeepONet. DeepONet consists of two subnetworks: the branch net and the trunk net. In quantum DeepONet, we replace these with QOrthoNN, which is composed of several quantum layers arranged sequentially. The addition of biases and nonlinear operations are performed on classical computers.
@@ 2.3 Quantum DeepONet
A DeepONet includes a branch net and trunk net, each with an equivalent number of output neuron, denoted by $p$. The branch and trunk nets can adopt arbitrary architectures, like fully connected neural network (FNN), convolutional neural network (CNN), recurrent neural network (RNN), and residual neural network (ResNet). A diagrammatic representation of DeepONet is illustrated in the center of Fig. 3. The branch network receives the input function evaluated at a discrete set of points $\{z_{1},z_{2},\ldots,z_{q}\}$, represented by $[v(z_{1}),v(z_{2}),\ldots,v(z_{q})]$. The trunk net is fed with the location $\xi$ at which the output function is evaluated, which can include both time and space coordinates. The outputs of the branch and trunk networks are denoted by $[b_{1}(v),b_{2}(v),\ldots,b_{p}(v)]$ and $[t_{1}(\xi),t_{2}(\xi),\ldots,t_{p}(\xi)]$. Thus, the final output of DeepONet is the sum of the dot product of the branch and trunk network outputs and a bias $b_{0}\in\mathbb{R}$, expressed as
@@ 2.4 Training quantum DeepONet
Hence, the primary speedup offered by our method originates from the evaluation phase, while the training phase maintains a complexity comparable to that of standard neural networks.
@@ 3 Ideal quantum simulation results
To demonstrate the efficacy of our method, we first use QOrthoNN to approximate certain functions (Section 3.1). Subsequently, we move to the application of quantum DeepONet on learning ODE and PDE problems, including the antiderivative operator (Section 3.2), advection equation (Section 3.3), and Burgers’ equation (Section 3.4). Finally, we test QPI-DeepONet using the antiderivative operator and Poisson’s equation (Section 3.5).
We implement the classical training by using the library DeepXDE [8]. After classical training on OrthoNN, we extract the weights and biases and construct a quantum version incorporating quantum layers, applying Qiskit [54] for quantum simulation. It is important to note that, in this section, we adopt an idealized scenario during quantum simulation. This approach excludes any quantum and statistical noise, aiming to assess the theoretical accuracy and performance of the quantum model. Table 3 presents the neural network hyperparameters, such as the learning rate and the number of iterations used in Adam optimization [55], along with the errors of different examples. The code of all examples is published on GitHub (https://github.com/lu-group/quantum-deeponet).
Table 3: The default parameters and test $L^{2}$ relative error for different examples of quantum DeepONet. For quantum DeepONet, the first number in the “Depth” column is the depth of the branch net, and the second number is the depth of the trunk net. The same for the “Activation” column. The “Error” column presents results for both classical training and ideal simulation, as they are identical for all examples. Here, “classical” denotes the use of classical DeepONet to benchmark against the quantum DeepONet. In classical DeepONet, we have utilized a smaller network depth to ensure that the number of trainable parameters remains comparable to those in the quantum version.
@@ 3.1 Function approximation
and approximate it using OrthoNN. We choose 80 points for training and 100 points for testing, where $x$ is uniformly sampled in $[-1,1]$. Specifically, for this example, we use the tanh activation function to circumvent the “dying ReLU” problem [56], which is particularly relevant here given the small width of the network.
For this function 1, we can achieve a small $L^{2}$ relative error of $0.149\%$ for the testing set after training classically. Following classical training, we construct QOrthoNN using the pyramidal quantum circuit, based on the classically training parameter. The ideal quantum simulation yields an error identical to classical training: $0.149\%$. Essentially, OrthoNN and QOrthoNN are the same neural network, differing only in their prediction methods—one is executed on a classical computer, while the other is run on a quantum simulator. The results for true function, OrthoNN, and QOrthoNN are plotted in Fig. 4A, where the three lines align closely with each other.
Figure 4: Quantum simulation result of function predictions. The black, red, and blue lines represent the reference solution, classical prediction of OrthoNN, and ideal quantum simulation result of QOrthoNN, respectively. (A) Results for $f(x)=1/(1+25x^{2})$. (B) Results for $f(x)=\sum_{k=1}^{4}\sin(kx)$.
@@ 3.2 Antiderivative operator
Next, we examine quantum DeepONet by starting with an antiderivate operator:
@@
where $k_{l}(x_{i},x_{j})=\exp\left(-\frac{d(x_{i},x_{j})^{2}}{2l^{2}}\right)$ denotes the radial basis function (RBF) kernel. In this context, $d(\cdot,\cdot)$ is the Euclidean distance between two points, and $l$ represents the length scale of the kernel, which modulates the smoothness of the generated function. Specifically, an increase of the value of $l$ leads to a smoother generated function. Therefore, we can adjust $l$ depending on our desired level of function’s complexity.
In this example, we explore two scenarios with different length scales: $l=1.0$ and $l=0.5$, corresponding to different size of the network during training. We achieved small errors of $0.49\%$ and $0.84\%$, respectively in these two scenarios for both the training of quantum DeepONet and ideal simulation (Table 3).
@@ 3.3 Advection Equation
$\frac{\partial u}{\partial t}+\frac{\partial u}{\partial x}=0,\quad x\in[0,1],\leavevmode\nobreak\ t\in[0,1],$
@@
$k(x_{i},x_{j})=\exp{\left(-\frac{2\sin^{2}{(\pi d(x_{i},x_{j})/p)}}{l^{2}}\right)}.$
Here, $p$ is the periodicity of the kernel and is set to $1$. We choose $l=1.5$ and derive the ground truth using the analytical solution $u(x,t)=u_{0}(x-t)$. For branch inputs $u_{0}(x)$, 20 sensors are uniformly placed (see one example in Fig. 5A left). Regarding the trunk inputs, we employ a grid of $50\times 50$ points, covering the range of $x$ and $t$. We implement the ResNet [57] architecture in both branch and trunk nets, which has a formulation of $\mathbf{x}^{\prime}=\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})+\mathbf{x}$ for each layer. This approach effectively mitigate the issue of gradient vanishing during training. The final test error of classical prediction reaches $2.25\%$. Ideal simulation of quantum DeepONet yields the same error: $2.25\%$. The quantum DeepONet, with $3081$ trainable parameters, matches the accuracy of the classical DeepONet, which achieves an error of $1.91\%$ using $3171$ trainable parameters (Table 3). Fig. 5A provides an example of illustrating the ground truth, predictions of quantum DeepONet, and the absolute error between them.
@@ 3.4 Burgers’ Equation
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},\quad x\in[0,1],\leavevmode\nobreak\ t\in[0,1],$
with initial condition $u_{0}(x)$ and periodic boundary condition, where $\nu=0.05$ is the viscosity. We aim to learn the mapping from $u_{0}(x)$ to the solution $u(x,t)$. Additionally, recognizing the periodic nature of the output function $u(x,t)$, instead of directly input $\xi$ for trunk net, we expand it to $[\xi,\cos(2\pi\xi),\sin(2\pi\xi),\cos(4\pi\xi),\sin(4\pi\xi)]$. Other neural network settings are the same as Section 3.3, except for the depth and width. With comparable number of trainable parameters ($2429$ for quantum DeepONet and $2260$ for classical DeepONet), the classical DeepONet achieves an error of $1.05\%$, whereas the quantum DeepONet sustains comparable performance, exhibiting a classical test relative error of $1.38\%$, and an ideal quantum simulation test error of $1.38\%$. An example of the ground truth and prediction of quantum DeepONet is shown in Fig. 5B.
@@ 3.5 Quantum physics-informed DeepONet
In this section, we further show that our quantum DeepONet can also be trained using physics-informed loss term (Eq. (4)), without labeled data. We choose the antiderivative equation in Eq. (5) for comparison with a data-driven case in Section 3.2. Additionally, 1D Poisson’s equation
@@
with zero Dirichlet boundary condition is also considered for demonstration. The branch inputs in both cases are the $v(x)$ in equations, which are generated by GRF with RBF kernel. To facilitate training and keep PDE residual within a reasonable range, in Poisson’s equation, we multiply the generated GRF with a factor of 10 and take the enlarged function as an input sample. The boundary condition is hard constrained using corresponding neural network architecture as mentioned in Section 2.5. Zero coordinate shift algorithm [58] is utilized to reduce GPU memory consumption and training time. During the training, the number of input samples is 10000 with batch size = 2000. Adam optimization is used with $2\times 10^{5}$ iteration. The PDE residual is evaluated at 100 uniformly distributed points in $[0,1]$. During the training, we employed PCA with the original dimension 100. The training results are shown in Table. 4. Ideal quantum simulation results also agree well with classical training results for all of these examples, shown “$L^{2}$ relative error” column.
We also conducted experiments without any dimension reduction techniques. For antiderivative with initial condition $l=1$, using branch and trunk net with a depth of 3 and width of 10, the test $L^{2}$ relative error is $4.05\%$. For comparison, using the same hyperparameters with PCA, which projects the original 100 dimensions down to 10, resulted in an error of $0.76\%$. We believe this difference is due to the critical dependency of QPI-DeepONet on the sampling of input sensors. As derivatives are taken with respect to the inputs, QPI-DeepONet is more sensitive to the input data. PCA enables us to incorporate more information within a limited input dimension. The limitations of current quantum devices compel us to use narrower neural networks, leading to sparse sampling of the branch input.
@@ 4 Effects of noise
Quantum noise is a major obstacle to the practicality of a quantum algorithm in the noisy intermediate-scale quantum (NISQ) era. It emerges from various sources, including the imperfect implementation of quantum operators, undesired environmental or qubit interactions, and erroneous state preparation or measurement. During the execution of a quantum circuit, the accumulated errors produced by the noise can destroy any information we intend to obtain. Meanwhile, even in a fault-tolerant scenario, the inaccuracy resulting from the finite number of measurements still affects the error level and complexity of the neural network, making it unavoidable to discuss the feasibility of our work on near-term quantum computers. Thus, in Sections 4.1 and 4.2, we first provide a theoretical analysis of the effects of finite-sampling noise on the single RBS gate and tomography outputs and a well-known noise channel, depolarizing noise, respectively. Then, we demonstrate our noisy simulation results of quantum DeepONet under both types of noise, as well as a more comprehensive noise model emulating a real IBM quantum computer in Section 4.3.
@@ 4.1 Finite-sampling noise in tomography
In the tomography step (Section 2.1.4), the probabilities ${\rm Pr}[0,\mathbf{e}_{j}]$ and ${\rm Pr}[1,\mathbf{e}_{j}]$, for $j\in\{1,...,r\}$, are estimated from the frequencies of measurement outcomes, where $n$ is the input vector dimension, $m$ is the output vector dimension, and $r=\max(m,n)$. This introduces an additional error in the estimation of output vector $\mathbf{y}$, caused by the finite number of measurements (shots). Let $q^{(0,j)}$ be the probability of measuring |$\ket{0,\mathbf{e}_{j}}$ in tomography layer, and $\hat{q}^{(0,j)}$ be the estimated value from $N_{shot}$ shots. To estimate the size of the finite-sampling error, we first calculate the standard deviation of $\hat{q}^{(0,j)}$.
@@
and $S^{(0,j)}=\sum_{k=1}^{N_{shot}}Z^{(0,j)}_{k}\sim\mathrm{Bin}\left(N_{shot},q^{(0,j)}\right)$. Thus, we have
@@
It follows that $\mathrm{E}\left[\hat{q}^{(0,j)}\right]=q^{(0,j)}$ and, asymptotically, the delta method yields
[…]
$\mathrm{Std}\left[y_{j}\right]$ $=\mathrm{Std}\left[{\rm sign}(y_{j})\left(2\sqrt{{\rm Pr}[0,\mathbf{e}_{j}]}-\frac{1}{\sqrt{r}}\right)\right]$
$=2\cdot\mathrm{Std}\left[\sqrt{\hat{q}^{(0,j)}}\right]\approx\frac{\sqrt{1-q^{(0,j)}}}{\sqrt{N_{shot}}}\propto\frac{1}{\sqrt{N_{shot}}},$ (6)
where $q^{(0,j)}\in[0,1]$. Similarly, defining $\hat{q}^{(1,j)}$ as the probability of measuring |$\ket{1,\mathbf{e}_{j}}$ in tomography layer and letting $\hat{q}^{(1,j)}$ be its estimate from $N_{shot}$ shots, we also have $\mathrm{Std}\left[y_{j}\right]\propto 1/N^{-0.5}_{shot}$ for $y_{j}<0$. In conclusion, the finite-sampling error on the estimation of output vector $\mathbf{y}\in\mathbb{R}^{r}$ is proportional to $N_{shot}^{-0.5}$ when $N_{shot}$ is large enough.